OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6) y=1/3x+7 y=-3x+3 y=1/3x+3 y=-3x+7

OpenStudy (anonymous):

since we have a point on the line we want, we need the slope. in order to be parallel, we must have the same slope as the given line. what's the slope?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

nope.. y = -3x + 12 what's the slope of this line? remember y=mx+b, where m is the slope.

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

ok good.

OpenStudy (anonymous):

so now you have slope, m=-3, and a point the line goes through, (-1, 6) we can put that in point-slope form: y - y1 = m(x - x1)

OpenStudy (anonymous):

once you have that, do some algebra and solve for y so you can make the appropriate choice... :)

OpenStudy (anonymous):

How do I put it in the Slope-point form?

OpenStudy (anonymous):

sorry, (x1, y1) is the point (-1, 6), and m is our slope. copy my drawing and fill in those boxes.